OpenStudy (dandandan):

T: R^3 --> R^3 such that T(1,1,1)=(2,0,1) T(0,1,0)=(-3,2-1) T(1,0,1)=(1,1,0) find T(2,-1,1)

OpenStudy (jamesj):

You want to write (2,-1,1) as a linear combination of the other three 'input' vectors that you are given: (2,-1,1) = a(1,1,1) + b(0,1,0) + c(1,0,1) Then T(2,-1,1) = aT(1,1,1) + bT(0,1,0) + cT(1,0,1) If you just look at this for a moment, you should be able to figure a,b,c out. But if you're really really stuck, solve the system of equations.

OpenStudy (dandandan):

hey, the values of a, b and c are not really necessary to be equal to (2,-1,1) right ?

OpenStudy (jamesj):

I don't quite understand what you're asking. If you're asking "(a,b,c) is not necessarily equal to (2,-1,1)" the answer is no, they are in fact not. But if you're asking, do we need to find the constants a,b,c such that (2,-1,1) = a(1,1,1) + b(0,1,0) + c(1,0,1) the answer is an definite yes.

OpenStudy (dandandan):

yeah, im talking about the first one, okay sir ! thanks a lot !!

OpenStudy (jamesj):

Hang on a second: are you sure these equations are correct? T(1,1,1)=(2,0,1) T(0,1,0)=(-3,2-1) T(1,0,1)=(1,1,0) In particular, I think one of the vectors on the left side is not correct.

OpenStudy (jamesj):

and that's because the three vectors: (1,1,1), (0,1,0) and (1,0,1) are not linearly independent and do not span R^3. In particular, there are no constants a,b,c such that: (2,-1,1) = a(1,1,1) + b(0,1,0) + c(1,0,1)

OpenStudy (dandandan):

okok ! wait sir, how can i know if a S spans R^3 ?

OpenStudy (dandandan):

coz i just know that if the determinant of S in non zero then it spans R^3

OpenStudy (jamesj):

Yes, if you took the determinant of the matrix comprising of these three vectors as columns or rows, it would be zero. In any case, double check the question.

OpenStudy (dandandan):

ok, so therefore if determinant is zero, is does not span?

OpenStudy (jamesj):

yes